# Invariance of deep image quality metrics to affine transformations

### v1

@@ Abstract
In this work, we evaluate state-of-the-art deep image quality metrics by assessing their invariance to affine transformations, specifically: rotation, translation, scaling, and changes in spectral illumination. We propose a methodology to assign invisibility thresholds for any perceptual metric. This methodology involves transforming the distance measured by an arbitrary metric to a common distance representation based on available subjectively rated databases. We psychophysically measure an absolute detection threshold in that common representation and express it in the physical units of each affine transform for each metric. By doing so, we allow the analyzed metrics to be directly comparable with actual human thresholds. We find that none of the state-of-the-art metrics shows human-like results under this strong test based on invisibility thresholds.
This means that tuning the models exclusively to predict the visibility of generic distortions may disregard other properties of human vision as for instance invariances or invisibility thresholds. The code is publicly available to test other metrics  [^1] https://github.com/Rietta5/InvarianceTestIQA.
@@ I Introduction
Deep architectures are the current state-of-the-art in predicting subjective image quality. These models, usually referred to as perceptual metrics, are often used as measures to optimize other models [1, 2]. This implies that assessing their performance can be critical in a wide range of applications. Usually, these models are evaluated (or even tuned) according to their ability to correlate with human opinion in databases including a wide range of generic distortions [3, 4]. In fundamental terms, this means trying to predict the visibility of generic distortions. However, focus on generic distortions may be a problem to take into account other relevant phenomena [5], and human vision may also be described in terms of perceptual constancies or invariances [6, 7, 8, 9, 10, 11].
@@
In fact, the spirit of the influential SSIM was focused on measuring changes of structure [13] to achieve invariances to irrelevant transformations [14]. Moreover, Wang and Simoncelli [12] decomposed generic distortions into structural and non-structural components so that the part not affecting the structure (e.g. affine transformations) could be processed, and weighted, differently. On the other hand, metrics with bio-inspired, explainable architecture [15, 16, 17, 18, 19, 20, 5, 4] work in multi-scale / multi-orientation domains where invariances can be introduced by means of appropriate poolings as in the scattering transformations [21]. This sort of poolings are thought to happen in the visual brain leading to invariances and texture metamers [22].
@@
In this work, we compare the ability of metrics to be invariant to affine transformations in the same way as humans are. In particular, we propose to evaluate the metrics from the point of view of human detection thresholds: by assessing the invariance of the metrics to image transformations that are irrelevant (or invisible) to human observers. For example, the classical literature on visual thresholds determines the intensity of certain affine transformations which is invisible for humans [24, 25, 26, 27]. The sizes of invisibility thresholds are related to the more general concept of invariance to transformations. By definition, transformations whose intensity is below the threshold are invisible to the observer. Then one can say that, in this region, the observer is invariant to the transform. Here we propose a methodology that allows us to assign specific invisibility thresholds per metric. This proposal uses transduction functions to a common internal representation based on a subjectively rated database and a psychophysically measured threshold in this representation. Then, we can (1) assess if the thresholds for the metrics are comparable to those found for human observers, and (2) assess if the sensitivities of a metric for the different distortions follow the same order as the human sensitivities (for instance humans are more sensible to rotation changes than to illumination changes). This proposed evaluation of invariance to distortions is a necessary complement to the conventional evaluation of the visibility of distortions because, as we will see, none of the studied metrics presents human-like thresholds nor sensitivities for all the transformations considered.
@@ II Proposed Methodology: Thresholds and Sensitivities
While invariance thresholds in human observers, $\theta_{\tau}^{\textrm{H}}$, are easy to understand and measure [24, 25, 26, 27], they are not obvious to define in artificial models of perceptual distance. The reason is simple: for the usual image quality models any non-zero image distortion leads to non-zero variations in the distance. In this situation where artificial distances are real-valued functions one should define a value, the threshold distance, $\mathcal{D}_{\tau}$, below which the difference between the images could be disregarded (or taken as zero). Once this threshold distance is available (eventually in a common scale for all metrics) one can translate this threshold to the axis that measures the distortion intensity, $\theta$, and hence obtain the threshold of the metric, $\theta_{\tau}^{M}$, in the same units that has been measured for humans, $\theta_{\tau}^{H}$. See an illustration of this concept in Fig. 1.
The diagram in Fig. 1 displays the relation between the physical description of the intensity of certain distortions in abscissas (the parameter $\theta$, e.g. the angle in a rotation transform), and the common internal description of the perceived distance in ordinates (the distance $\mathcal{D}$, e.g. the normalized Mean Opinion Score units explained below). This relation between the physical description of the intensity and common perceptual distance is what we call transduction function. The example displays three transduction functions, $\mathcal{D}_{M}=g_{M}(\theta)$, with $M=1,2,3$, in blue, red, and green, for the three corresponding metrics. Transduction functions are monotonic: the application of a progressively increasing distortion along the abscissas leads to monotonic increments in distances in ordinates. The threshold distance, $\mathcal{D}_{\tau}$, in the internal representation is plotted in orange in Fig. 1. This threshold may be uncertain (as represented by the central value, solid line, and the quartile limits, in dashed lines), but the empirical transduction functions can be used to put this internal threshold back in the axis that describes the distortion intensity: $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. In this way, one can check if the actual invisibility threshold measured for humans (black line) is consistent with the threshold interval deduced for each metric. In our illustration, the metric 2 (red line) is the only one compatible with human behavior.
@@
Figure 1: Illustration of the methodology: comparison of metric vs human thresholds. For each distortion, we evaluate three different metrics (continuous curves in blue, red, and green) in common units of subjective distance ($\mathcal{D}$, the normalized Mean Opinion Score). Given $N$ competing metrics, these curves are the transduction functions that map the physical description of the distortion intensity, $\theta$ into a common scale of subjective distance, $\mathcal{D}_{M}=g_{M}(\theta)$, with $M=1,\ldots,3$. The yellow lines show the human threshold in the internal distance representation, $\mathcal{D}_{\tau}$, experimentally measured in this work with a psychometric function. We assign the metric thresholds from the generic human threshold using the inverse of the transduction functions for the different metrics: $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. Given the fact that $\mathcal{D}_{\tau}$ has an experimental uncertainty, the inverse gives us an interval - dotted lines- on the x-axis where the human threshold should fall for each metric and this particular distortion. For any number $N$ of metrics we propose to compare the metric thresholds, $\theta_{\tau}^{M}$, for $M=1,\ldots,N$, with the actual human threshold in intensity units, $\theta_{\tau}^{H}$. In this case, $\theta_{\tau}^{H}$ (the black line), falls within the red interval. Therefore, the metric 2 is the one with the human-like behavior.
@@
Using the above, on the one hand, we propose to evaluate the alignment between the metric models and human observers by comparing $\theta_{\tau}^{M}$ versus $\theta_{\tau}^{H}$. This is a strict comparison that only depends on the experimental uncertainty of the thresholds. On the other hand, we can define an alternative (less strict) comparison by considering the order among the sensitivities for the different distortions in humans and models. In particular, one can define the sensitivity of a metric for a distortion as the variation of the transduction function in terms of the energy (or Mean Squared Error) of the distortion introduced by the transform. The human sensitivity in detection is classically defined to be proportional the inverse of the energy required to see the distortion [28].
@@
    The transduction function (red, green, and blue curves in the illustration). For the $n$-th metric, $d_{M}(i,i^{\prime})$, the relation between the physical description of the image transform and a common metric-independent distance domain has two components:
@@
      A metric equalization function transforming the previous (non-scaled) distance values into the common scale of the internal distance representation (what we called normalized DMOS units in the illustration). Here we propose to use auxiliary empirical data (e.g. certain subjectively rated databases) to scale the range of the different metrics: $\mathcal{D}_{M}=f_{M}(d_{M})$. This makes different $\mathcal{D}_{M}$ comparable.
@@
      In the units of the own metric, $d_{\tau}^{M}=f_{M}^{-1}(\mathcal{D}_{\tau})$. This value is particularly interesting, as it indicates a variation in the distance of each metric for which humans see no difference between the original and the distorted image. Distortions leading to distances below this value are invisible to humans and hence should be neglected.
@@ II-B Human thresholds
In this section, we measure the human thresholds from two different points of view. First, we use distorted images from a subjectively rated database to measure from the internal common representation. Then, we measure it from the internal physical representation using natural stimuli.
@@ B.1. Human thresholds in the common internal representation
The value of $\mathcal{D}_{\tau}$ could be roughly estimated by visual inspection of the images presented in Appendix A: images with low values of normalized DMOS (below 0.3) cannot be discriminated from the original, while images with big normalized DMOS (above 0.6) are clearly distinct from the original. However, the more accurate estimation of such a threshold is obtained from the psychometric function in a constant stimulus experiment [29] applied to other fields within computer vision, such as [36, 37, 38, 39, 40]
@@ B.2. Human threshold in physical units
Humans are not equally sensitive to all transformations or to different degrees of distortion, $\theta$. In fact, there are certain thresholds at which these distortions are not perceptible, $\theta_{\tau}^{H}$, known as the invisibility threshold. Initially, we were going to use the thresholds from the classical literature [24, 25, 26] but, considering that these thresholds are measured with synthetic stimuli and not with natural images, we decided to do the experiments ourselves. To derive the new thresholds with human data, we followed the psychometric function methodology, following the Equation 4. Contrary to the literature, these experiments have been done with natural images, to use the same type of images that the models use. In particular, images from ImageNet have been used to find the invisibility thresholds in a more realistic context. Regarding the geometric affine transformations (Section III-A), the procedure follows the classic psychometric function methodology (Section II-B). The results of the new measurements on humans with natural stimuli can be seen in Figure 4.
@@
Figure 4: Experimental (human) psychometric functions measured with natural images. According to them, the human thresholds for the geometric affine transformations are 0.23 degrees for translation, 3.6 degrees for rotation, and 1.026 for scale.
Because when changing the spectral illuminant we move in a 2D space, illuminant transformations are treated slightly differently. On the one hand, the MacAdam ellipses [27] don’t contain any ellipse centered in the true white, coordinates $(1/3,1/3)$ in the spectral locus, but it is the starting point of the employed illuminants. This can be solved by fitting a new ellipse at this point considering the closest existing ellipses. On the other hand, experimentally evaluating a new ellipse employing natural images would imply, in our case, fitting 20 psychometric functions. Instead, we only obtain the psychometric functions corresponding to the major and minor axes of the ellipse: Yellow-Blue (YB) and Red-Green (RG) chromatic directions respectively. For a more accurate fit, two intermediate hues have also been measured. The fitted psychometric functions for these 4 chromatic directions are shown in Figure 5.
@@
Once these points are measured, we can fit a new experimental ellipse that has been measured only with natural images. Both ellipses, the fitted and the experimental ones, have been obtained with the calculations from [41], and can be seen in Figure 6.
Figure 6: New fitted ellipses. In purple, the new ellipse centered on the absolute white interpolated from the closest ellipses (in green). In red, the experimental ellipse fitted with our own measured data. The highlighted axes are the interpolated ellipse’s major and minor axes, corresponding to hues 4 and 8 (orange and cyan).
@@
Translation 0.024 degrees [24] 0.23 degrees
Rotation 3 degrees [25] 3.6 degrees
Scale 1.03 scale factor [26] 1.026 scale factor
Color discriminant MacAdam ellipse [27] Experimental ellipse (Fig. 6)
From Table I we see that the rotation and scale thresholds are practically equivalent, but this is not the case for the translation threshold. This can be attributed to the fact that experiments with synthetic stimuli can be considered as being close to a hyperacuity setting. That is, they are based on an experiment in which images are shown in successive comparisons rather than in simultaneous comparisons, as in our case. Our experiments employ natural images and do not enforce the hyperacuity setting. This results in a considerably higher detection threshold for this transformation. Another more recent work we could consider in this area are the thresholds measured in 2016 with gabor filters [42]. In this case, the human rotation threshold was considered to be 2.7 degrees of rotation and the human translation threshold the equivalent of 0.12 degrees of translation. These measurements are quite consistent with ours. Taking into account that the methodology of obtaining the threshold is the same as in our case, the difference in the thresholds is given exclusively by the type of stimulus used. For this reason, the detection threshold for translation from other literature will not be considered in the results because of differences in the procedure for obtaining it and its interpretation in the case of classical literature; and for the type of the stimulus used in the recent literature.

### v2

@@ Abstract
In this work, we evaluate state-of-the-art deep image quality metrics by assessing their invariance to affine transformations, specifically: rotation, translation, scaling, and changes in spectral illumination. Here invariance of a metric refers to the fact that certain distances should be neglected (considered to be zero) if their values are below a threshold. This is what we call invisibility threshold of a metric. We propose a methodology to assign such invisibility thresholds for any perceptual metric. This methodology involves transformations to a distance space common to any metric, and psychophysical measurements of thresholds in this common space. By doing so, we allow the analyzed metrics to be directly comparable with actual human thresholds. We find that none of the state-of-the-art metrics shows human-like results under this strong test based on invisibility thresholds.
This means that tuning the models exclusively to predict the visibility of generic distortions may disregard other properties of human vision as for instance invariances or invisibility thresholds. The data and code is publicly available to test other metrics [^1] https://github.com/Rietta5/InvarianceTestIQA.
@@ I Introduction
Deep architectures are the current state-of-the-art in predicting subjective image quality. These models, usually referred to as perceptual metrics, are often used as measures to optimize other image processing tasks [1, 2] so assessing their performance is critical. Usually, these models are evaluated (or even tuned) according to their ability to correlate with human opinion in databases including a wide range of generic distortions [3, 4]. In fundamental terms, this means trying to predict the visibility of generic distortions. However, focus on generic distortions may be a problem to take into account other relevant phenomena [5], and human vision may also be described in terms of perceptual constancies or invariances [6, 7, 8, 9, 10, 11].
@@
In fact, the spirit of the influential SSIM was focused on measuring changes of structure [13] to achieve invariances to irrelevant transformations [14]. Moreover, Wang and Simoncelli [12] decomposed generic distortions into structural and non-structural components so that the part not affecting the structure (e.g. affine transformations) could be processed, and weighted, differently. On the other hand, metrics with bio-inspired, explainable architecture [15, 16, 17, 18, 19, 20, 5, 4] work in multi-scale / multi-orientation representations where invariances can be introduced by means of appropriate poolings as in the representations [21]. This sort of poolings are thought to happen in the visual brain leading to invariances and texture metamers [22].
@@
In this work, we compare the ability of metrics to be invariant to affine transformations in the same way as humans are. In particular, we propose to evaluate the metrics from the point of view of human detection thresholds: by assessing the invariance of the metrics to image transformations that are irrelevant (or invisible) to human observers. For example, the classical literature on visual thresholds determines the intensity of certain affine transformations which is invisible for humans [24, 25, 26, 27]. The sizes of invisibility thresholds are related to the more general concept of invariance to transformations. By definition, transformations whose intensity is below the threshold are invisible to the observer. Then one can say that, in this region, the observer is invariant to the transform. Here we propose a methodology that allows us to assign specific invisibility thresholds per metric. This proposal uses transduction functions to a common internal representation based on a subjectively rated database and a psychophysically measured threshold in this representation.
Then, we can (1) assess if the thresholds for the metrics are comparable to those found for human observers, and (2) assess if the sensitivities of a metric for the different distortions follow the same order as the human sensitivities (for instance humans are more sensible to rotation changes than to illumination changes). This proposed evaluation of invariance to distortions is a necessary complement to the conventional evaluation of the visibility of distortions because, as we will see, none of the studied metrics presents human-like thresholds nor sensitivities for all the transformations considered.
@@ II Proposed Methodology: Thresholds and Sensitivities
While invariance thresholds in human observers, $\theta_{\tau}^{\textrm{H}}$, are easy to understand and measure [24, 25, 26, 27], they are not obvious to define in artificial models of perceptual distance. The reason is simple: for the usual image quality models any non-zero image distortion leads to non-zero variations in the distance. In this situation where artificial distances are real-valued functions one should define a value, the threshold distance, $\mathcal{D}_{\tau}$, below which the difference between the images could be disregarded (or taken as zero). Once this threshold distance is available (eventually in a common scale for all metrics) one can translate this threshold to the axis that measures the distortion intensity, $\theta$, and hence obtain the threshold of the metric, $\theta_{\tau}^{M}$, in the same units that has been measured for humans, $\theta_{\tau}^{H}$. See an illustration of this concept in Figure 1.
@@
Figure 1: Illustration of the methodology: comparison of metric vs human thresholds. For each distortion, parametrized with $\theta$, we evaluate three different metrics (continuous curves in blue, red, and green) in common units of subjective distance ($\mathcal{D}$, the normalized Mean Opinion Score). Given $N$ competing metrics, these curves are the transduction functions that map the physical description of the distortion intensity, $\theta$ into a common scale of subjective distance, $\mathcal{D}_{M}=g_{M}(\theta)$, with $M=1,\ldots,3$. The yellow lines show the human threshold in the internal distance representation, $\mathcal{D}_{\tau}$, experimentally measured in this work with a psychometric function. We assign the metric thresholds from the generic human threshold using the inverse of the transduction functions for the different metrics: $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. Given the fact that $\mathcal{D}_{\tau}$ has an experimental uncertainty, the inverse gives us an interval - dotted lines- on the x-axis where the human threshold should fall for each metric and this particular distortion. For any number $N$ of metrics we propose to compare the metric thresholds, $\theta_{\tau}^{M}$, for $M=1,\ldots,N$, with the actual human threshold in intensity units, $\theta_{\tau}^{H}$. In this case, $\theta_{\tau}^{H}$ (the black line), falls within the red interval. Therefore, the metric 2 is the one with the human-like behavior.
The diagram in Figure 1 displays the relation between the physical description of the intensity of certain distortions in abscissas (the parameter $\theta$, e.g. the angle in a rotation transform), and the common internal description of the perceived distance in ordinates (the distance $\mathcal{D}$, e.g. the normalized Mean Opinion Score units explained below). This relation between the physical description of the intensity and common perceptual distance is what we call transduction function. The example displays three transduction functions, $\mathcal{D}_{M}=g_{M}(\theta)$, with $M=1,2,3$, in blue, red, and green, for the three corresponding metrics. Transduction functions are monotonic: the application of a progressively increasing distortion along the abscissas leads to monotonic increments in distances in ordinates. The threshold distance, $\mathcal{D}_{\tau}$, in the internal representation is plotted in orange in Figure 1. This threshold may be uncertain (as represented by the central value, solid line, and the quartile limits, in dashed lines), but the empirical transduction functions can be used to put this internal threshold back in the axis that describes the distortion intensity: $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. In this way, one can check if the actual invisibility threshold measured for humans (black line) is consistent with the threshold interval deduced for each metric. In our illustration, the metric 2 (red lines) is the only one compatible with human behavior.
@@
Using the above, on the one hand, we propose to evaluate the alignment between the metric models and human observers by comparing $\theta_{\tau}^{M}$ versus $\theta_{\tau}^{H}$. This is a strict comparison that only depends on the experimental uncertainty of the thresholds. On the other hand, we can define an alternative (less strict) comparison by considering the order among the sensitivities for the different distortions in humans and models. In particular, one can define the sensitivity of a metric for a distortion as the variation of the transduction function in terms of the energy (or Mean Squared Error) of the distortion introduced by the transform [28]. The human sensitivity in detection is classically defined to be proportional the inverse of the energy required to see the distortion [29].
@@
    The transduction function (red, green, and blue curves in the illustration). For the $M$-th metric the relation between the physical description of the image transform and a common metric-independent distance domain has two components:
@@
      A metric equalization function transforming the previous (non-scaled) distance values into the common scale of the internal distance representation (what we called normalized MOS units in the illustration). Here we propose to use auxiliary empirical data (e.g. certain subjectively rated databases) to scale the range of the different metrics: $\mathcal{D}_{M}=f_{M}(d_{M})$. This makes the different $\mathcal{D}_{M}$ comparable.
@@
      In the units of the metric, $d_{\tau}^{M}=f_{M}^{-1}(\mathcal{D}_{\tau})$. This value is particularly interesting, as it indicates a variation in the distance of each metric for which humans see no difference between the original and the distorted image. Distortions leading to distances below this value are invisible to humans and hence should be neglected.
@@ II-B Human thresholds
In this section, we measure two kinds of human thresholds. First, we use distorted images from a subjectively rated database to measure from the internal common representation, $\mathcal{D}_{\tau}$. Then, we measure the thresholds in physichal units, $\theta_{\tau}$, using natural stimuli.
@@ B.1. Human thresholds in the common internal representation
The value of $\mathcal{D}_{\tau}$ could be roughly estimated by visual inspection of the images presented in Appendix A: images with low values of normalized DMOS (below 0.3) cannot be discriminated from the original, while images with big normalized DMOS (above 0.6) are clearly distinct from the original. However, the more accurate estimation of such a threshold is obtained from the psychometric function in a constant stimulus experiment [30] applied to other fields within computer vision and image processing [36, 37, 38, 39, 40].
@@ B.2. Human threshold in physical units
In principle, one could take the thresholds $\theta_{\tau}^{H}$ from the classical literature [24, 25, 26, 27] but, considering the critical effect of the stimuli used in the experiments and the fact that classical thresholds were measured with synthetic stimuli, we decided to measure these thresholds for natural images to ensure consistency with the kind of images used by the metrics. We measure the thresholds as in Section II-B.1., according to the constant stimuli setting to derive the classical psychometric function [30]. In particular, images from ImageNet have been used to find the invisibility thresholds in a more realistic context.
For each transform, this experimental procedure requires the variation of the stimuli in the parameter space $\theta$, which is trivially one-dimensional for scale, translation, rotation, and intrinsically two-dimensional for chromatic shifts (despite they coming from arbitrary variations of a higher dimensional object such as the spectrum of the illuminant [41]).
On the one hand, regarding the geometrical transforms with one-dimensional parameter, the results of the thresholds obtained from the psychometric curves are shown in Figure 4.
@@
Figure 4: Experimental (human) psychometric functions measured with natural images. According to them, the human thresholds for the geometric affine transformations are 0.23 $\pm$ 0.10 degrees for translation, 3.6 ±$\pm 1.5$ degrees for rotation, and 1.03 ±$\pm 0.02$ for scale.
On the other hand, now we address the results of the color thresholds which are two-dimensional in nature. As a convenient summary, here we describe the chromatic content of natural images using a single two-dimensional point in the so-called chromatic diagram [41]. In particular, we will use the chromatic coordinates of the spectrum of the illuminant assumed to be the illumination source of the scene. Moreover, we will describe the color-shifts introduced to check the invariance of metrics by the new chromatic coordinates of the new spectrum applied to the scene (as affine transform on the image vector as in Eq. 8 in Section III-A). Here we assume that the natural images in the databases are, in principle, illuminated by a generic white spectrum. Therefore, we describe the chromaticity of these images using the location of that generic white spectrum in the CIE xy diagram, i.e. in the 2d-point $(1/3,1/3)$ [41].
The two-dimensional nature of chromatic shifts implies that invisibility regions are not intervals $[0,\theta_{\tau}^{H}]$, but two-dimensional shapes (e.g. discrimination ellipses [27]). In order to determine the shape of such discrimination region around a certain central chromatic location, e.g. the aforementioned white point, one could measure one-dimensional intervals moving the images in different chromatic directions away from the selected center. For instance, Figure 5 represents the white point $(1/3,1/3)$, generic illuminant, and four linear paths indicating color shifts away from this center (radial gray lines). Measurements along radial lines can be done with the constant-stimulus + psychometric function paradigm mentioned above. In our case, given the ellipsoidal shape of the classical chromatic discrimination thresholds [27] (also shown in Figure 5), measuring only in four directions is enough for a trustable fit of an ellipse [42]. The fitted psychometric functions for these four chromatic directions are shown in Figure 6. The thresholds in departure in chromatic coordinates from these psychometric functions determine four two-dimensional experimental thresholds that allow to fit the ellipse in pink in Figure 5.
Obtaining and assessing the result found (pink ellipse in Figure 5), and also the assessment of ellipses obtained from the metrics, depends on the measuring RMSE distances between ellipses. This error is relevant in the fitting process [42], where theoretical points are compared with the experimental points. Additionally, assessment implies the comparison of the predicted ellipsoid with the corresponding region that could be found in classical literature [27, 43]. As the selected reasonable white was not explicitly measured in the classical literature, we fitted a new ellipse at this point considering the closest existing ellipses (in green in Figure 5).
Figure 5: New fitted ellipses. In purple, the new ellipse centered on the absolute white interpolated from the closest ellipses (in green). In red, the experimental ellipse fitted with our own measured data. The highlighted axes are the interpolated ellipse’s major and minor axes and two intermediate hues.
Interestingly the ellipse we got by assessing the detection thresholds with color-shifted natural images is quite consistent with previous results that used synthetic stimuli and different experimental techniques [27, 43]. On the one hand, note the similarity in size and orientation between the pink and the purple ellipses. Note that the ellipse is wider in the yellow-blue (YB) direction than in the red-green (RG) direction. The authors in [43] reported that their threshold method with synthetic textured stimuli led to ellipses ×$\times 4.5$ than MacAdam ellipses [27] measured with the variance of color matching paradigms. As Fig 5 shows the MacAdam ellipses with a ×$\times 5$ visualization factor [41], our experimentally fitted pink ellipse obtained with thresholds from natural images is very consistent with [43].
@@
Translation 0.024 degrees [24] 0.23 $\pm$ 0.10 degrees
Rotation 3 degrees [25] 3.6 $\pm$ 1.5 degrees
Scale 1.03 scale factor [26] 1.03 $\pm$ 0.02 scale factor
Color discriminant MacAdam ellipse [27] Experimental ellipse (Figure 5)
From Table I we see that the rotation and scale thresholds are practically equivalent, but this is not the case for the translation threshold. This can be attributed to the fact their experiments with synthetic stimuli [24] can be considered as being close to a hyperacuity setting, (which returns very low thresholds) [44]. Our experiments employ natural images and do not enforce the hyperacuity setting. This results in a considerably higher detection threshold for this transformation. Recent work we could consider is the thresholds measured with gabor filters [45]. In this case, the human rotation threshold was considered to be 2.7 degrees of rotation and the human translation threshold the equivalent of 0.12 degrees of translation. These measurements are quite consistent with ours. Taking into account that the methodology of obtaining the threshold is the same as in our case, the difference in the thresholds is given exclusively by the type of stimulus used. For this reason, the detection threshold for translation from other literature will not be considered in the results because of differences in the procedure for obtaining it and its interpretation in the case of classical literature; and for the type of stimulus used in the recent literature.